When looking for a job, a person sends her application to 25 companies.

The probability that a company will answer is 0.2 and it is assumed that these answers are independent. 1) What is the probability, rounded to the nearest hundredth, that the person will receive at least one answer?
1
Expert's answer
2020-12-24T16:08:18-0500

We are dealing with the binomial distribution with parameters 25 and 0.2. We denote by "X" the respective variable. The probability is the following: "P(X\geq1)=1-P(X<1)=1-P(X=0)=1-(0.8)^{25}\approx1.00"

The latter is rounded to the nearest hundredth.